OpenStudy (valiyuh):

Find the equation perpendicular to 2y-4x=8 and passes the point (8,2) . Can someone explain how to do it please? -Geometry

OpenStudy (anonymous):

Change the given equation to slope intercept form, y=mx+b That will help you determine the slope of the given line

OpenStudy (anonymous):

What did you get

OpenStudy (valiyuh):

2y=4x+b ?

OpenStudy (anonymous):

close 2y = 4x + 8 No divide everything by 2 so you get y= mx+b

OpenStudy (anonymous):

Now

OpenStudy (anonymous):

did u get y=2x+4

OpenStudy (valiyuh):

yes

OpenStudy (anonymous):

You'll see that the slope of the given line is 2. To have a perpendicular line, the slope is the opposite reciprocal.....-1/2

OpenStudy (anonymous):

Using the slope intercept formula and the point given (8,2) plug in and find b.

OpenStudy (anonymous):

2=-1/2(8)+b

OpenStudy (anonymous):

Are you still there

OpenStudy (anonymous):

2=-4+b 6=b Now you can give the equation of a line perpendicular to the given line at the point (8,2). y=-1/2x+6

OpenStudy (valiyuh):

oh i see